First Assignment (due Thursday January 17) 1. Evaluate the following expression: 9+4X2/ (Y+1)-3Z where X 3, Y-5, Z-2 2. Solve the following equation for X where s 3, X-8 and z1. Invent a study with a psychologically relevant variable, such as compliance, aggression cooperation, happiness, etc. as the dependent variable. (In other words your dependent variable should not being something like course grade of blood pressure level, even if you independent variable is psychologically relevant). For example, you may want to see whether doctors or nurses are more skilled at producing compliance with patients. Make sure that your dependent variable is at least at the ordinal scale of measurement. 3. a. Describe an independent variable/dependent variable study b. What is your independent variable? c. What are the levels (groups) of your independent variable?
d. What is your dependent variable? Define it in general terms. e. Define your dependent variable operationally (i.e. tell how you would actually collect your data in terms that are so specific that any fool research assistant could follow them). Try not to use a rating scale to operationalize your variable, but if you do use a rating scale, then avoid a single rating number that directly asks to rate your degree of stress or happiness, etc. Instead, give some examples of specific questions that less directly get at what you are trying to measure. f. How well does the operational definition fit the general definition? Explain. g. What is the scale of of the dependent variahle (level of measurement)? Support your h. Is the dependent variable discrete or continuous? Support your answer. 氷! i. Is your design a true experiment or a pseudo experiment? Support your answer
Answer #1

1:

9+4X^{2}/(Y+1)-3Z=9+4\cdot (3^{2})/(5+1)-3\cdot 2=9+4\cdot (9)/(5+1)-3\cdot 2

=9+4\cdot (9)/(6)-6=9+6-6=9

Answer: 9

2:

z=\frac{X-\bar{X}}{s}=\frac{X-8}{3}

-1=\frac{X-8}{3}

X=-3+8=5

Answer: 5